### v1

# Robust Temporal-Invariant Learning in Multimodal Disentanglement

Abstract
Multimodal sentiment recognition aims to learn representations from different modalities to identify human emotions. However, previous works does not suppresses the frame-level redundancy inherent in continuous time series, resulting in incomplete modality representations with noise. To address this issue, we propose the Temporal-invariant learning, which minimizes the distributional differences between time steps to effectively capture smoother time series patterns, thereby enhancing the quality of the representations and robustness of the model. To fully exploit the rich semantic information in textual knowledge, we propose a Text-Driven Fusion Module (TDFM). To guide cross-modal interactions, TDFM evaluates the correlations between different modality through modality-invariant representations. Furthermore, we introduce a modality discriminator to disentangle modality-invariant and modality-specific subspaces. Experimental results on two public datasets demonstrate the superiority of our model.  [^1] Our code is available at https://github.com/X-G-Y/RTIL
Index Terms:
Multimodal Fusion, Temporal-Invariant Learning, Multimodal Disentanglement, Sentiment Analysis.
I Introduction
Figure 1: The overall structure of our proposed RTIL. In the feature extraction module, we begin by enriching the low-level features through the Transformers Encoders to obtain enhanced representations. The three processed modality embeddings are fed into shared and private encoders to extract the respective representations subsequently in the respresentation learning module. We use the consistency loss to constrain the modality-invariant subspace. Meanwhile, we separate the modality-specific subspace and the modality-invariant subspace by means of adversarial learning. Furthermore, the video features are constrained to learn the temporal-invariant representation. Eventually, the modality-specific features are fused in a text-driven manner within the fusion module, gated by the modality-invariant features.
Multimodal sentiment analysis (MSA) has become an active area of research with critical applications across various fields, such as human-computer interaction [1], social media analysis [2], and affective computing [3].
MSA typically involves data from video, speech, and text. Each of these modalities offers unique and complementary information that is crucial to the sentiment analysis. For instance, text data carries rich semantic content that directly conveys the speaker’s emotions and intentions. In contrast, video data provides valuable non-verbal cues, such as facial expressions and body language, which are essential for understanding the full context of the interaction.
Among the three modalities, the text modality stands out as the most dominant in MSA tasks [4]. Part of this advantage is due to the fact that text information is typically presented in a structured format, allowing for more precise expression of emotions and intentions. Another aspect is the continuous advancement of natural language processing techniques, which enable the accurate capture of emotional cues within text. Additionally, text data is more stable compared to audio and visual data, making it less susceptible to external factors.
In the context of multimodal fusion, the text modality complements other modalities effectively.Text can provide contextual information for audio and visual modalities, helping the model better understand the underlying emotional implications of speech or facial expressions. This complementarity between modalities significantly enhances the overall performance of the model, enabling multimodal sentiment analysis to capture more comprehensive and nuanced emotional information.
Compared to the semantically rich textual information, video data contains a significant amount of redundancy and noise. This redundancy stems from the high frame rate of video, where consecutive frames differ slightly, leading to repeated information that does not necessarily enhance emotion understanding. Additionally, noise in video data, such as changes in lighting conditions and background movements, further complicate the extraction of relevant emotional cues.
Based on the above observations, we propose temporal-invariant learning. Different from traditional complex preprocessing methods such as frame selection and noise reduction, temporal-invariant learning directly captures global information within video data at the feature level by minimizing the distributional differences between time steps. This approach effectively filters out the redundant and noisy information, allowing the model to capture the global information of the time series.
To address modality heterogeneity, we employ adversarial learning to train private and shared encoders to disentangle modality-specific and modality-invariant representations. Specifically, we utilize a spherical modality discriminative loss to enhance intra-class compactness and inter-class discrepancy for the hidden representations and parameters of the modality discriminator within a hyper-sphere [7].
Furthermore, we facilitate interactions between modalities with a focus on the text modality. To fully leverage the learned high-level shared representations, we propose an adaptive fusion mechanism that dynamically evaluates the correlations between modalities during the fusion process.
The main contributions can be summarised as follows: (1) We propose a novel multimodal representations learning model named RTIL (Robust Temporal-Invariant Learning), which leverages adversarial learning to separate modality-invariant representations and modality-specific representations, and adaptively fuses these representations. (2) Through temporal-invariant learning, redundant information and noise associated with temporal changes are effectively reduced, enhancing the model’s ability to maintain consistent global information regardless of variations in time steps. (3) We performed a series of experiments on two datasets, showing that the proposed RTIL outperforms state-of-the-art methods.
II PROPOSED MODEL
The overall architecture of RTIL is depicted in Fig. 1, consisting of the Feature Extraction Module, Representation Learning Module, and Text-Driven Fusion Module (TDFM). Further details are provided in the following subsections.
II-A Feature Extraction Module
For video and audio modalities, we use Transformer Encoders to capture long-range dependencies. For language modality, we feed the input text into RoBERTa to enhance the language representations. The outputs of each modality are denoted as $H_{i}$, where $i\in\{a,v,t\}$.
II-B Representation Learning Module
Modality-Invariant and Modality-Specific Representations Learning. RTIL leverages a shared encoder to capture invariant representations of different modalities, effectively reducing the heterogeneity gap. Additionally, to learn the specific representations, we utilize three different private encoders, mapping modalitiy embeddings to the modality-specific subspaces. The invariant representations $I_{i}$ and specific representations $S_{i}$ are denoted as:
$I_{i}=E_{I}(H_{i},\theta^{I}),S_{i}=E_{S}(H_{i},\theta^{S})$ (1)
where shared encoder $E_{I}$ shares the parameters $\theta^{I}$ and private encoders $E_{S}$ assign separate parameterare $\theta^{S}$ for each modality. To align the different modalities representations in the invariant subspace, we apply the consistency loss to disentangled representation learning. We use the Central Moment Discrepancy (CMD) [5] to measure the difference between two modalities.
Furthermore, the consistency loss can be calculated as:
$\mathcal{L}_{con}=\frac{1}{3}\sum_{m_{1},m_{2}\in\{a,v,t\}}CMD(I_{m_{1}},I_{m_{2}})$ (2)
Adversarial Learning. Inspired by the previous work [7], we introduce a modality discriminator to encourage the shared and private encoders to produce distinct representations. The invariant and specific representations are fed into the discriminator as input after passing through gradient reversal layers [8], then the discriminator predicts the modality from which the representation originates:
$\mathcal{D}({h_{i}},\theta_{\mathcal{D}})=softmax(\mathbf{W}_{\mathcal{D}}^{T}\cdot Linear({h_{i}}))$ (3)
where $h_{i}\in\{I_{i},S_{i}\}$ and $\mathbf{W}_{\mathcal{D}}$ is a learnable parameter matrix. We apply the additive angular margin loss [9] to enhance the intra-class compactness and inter-class discrepancy for the modality discriminator:
$\mathcal{L}_{am}=-log\frac{e^{\alpha\cdot\cos(\theta_{y_{m}}+\tau)}}{e^{\alpha\cdot\cos(\theta_{y_{m}}+\tau)}+\sum_{m=1,m\neq y_{m}}^{M}e^{\alpha\cdot\cos(\theta_{m})}}$ (4)
where $\hat{h}=Linear({h_{i}})$, $\theta_{y_{m}}=arccos(\mathbf{W}_{y_{m}}^{T}\cdot\hat{h_{i}})$ and $\theta_{m}=\arccos(\mathbf{W}_{m}^{T}\cdot\hat{h_{i}})$. $y_{m}$ denotes the ground-truth modality label. $\mathbf{W}_{y_{m}}$ denotes the $y_{m}$-th column of the weight matrix $\mathbf{W}_{D}$ and $\mathbf{W}_{m}$ denotes the $m$-th column of $\mathbf{W}_{D}$.
The domain loss can be calculated by combining the invariant and specific adversarial loss:
$\mathcal{L}_{dom}=\frac{1}{n}\sum_{i=1}^{n}\sum_{m\in\{a,t,v\}}(\mathcal{L}_{am}(I_{m},y_{m})+\mathcal{L}_{am}(S_{m},y_{m}))$ (5)
Figure 2: The details of Temporal-Invariant Learning.
Temporal-Invariant Learning. To further reinforce the temporal consistency of representations, we incorporate the concept of temporal-invariant learning.
Temporal-invariant learning aims to maintain features stable across time steps, ensuring that the learned representations are resilient by temporal variations. Specifically, we introduce the Jensen-Shannon [10] divergence (JSD) to constrain the video frames of a multivariate Gaussian distribution over the time steps as illustrated in Fig. 2.
The JSD is defined as:
$\text{JSD}(P\parallel Q)$ $=\frac{1}{2}\sum_{i}P(i)\log\left(\frac{P(i)}{M(i)}\right)$
$\quad+\frac{1}{2}\sum_{i}Q(i)\log\left(\frac{Q(i)}{M(i)}\right)$ (6)
where $M=\frac{1}{2}(P+Q)$ represents the average distribution of distribution $P$ and $Q$.
In video sequences, objects typically do not undergo significant changes between consecutive frames, resulting in a large amount of redundant information. Minimizing the distance between adjacent frames effectively reduces redundant information and thereby enhancing the stability and robustness of video representations. Based on this concept, our proposed temporal-invariance loss can be calculated as:
$\mathcal{L}_{ti}=\frac{1}{n-1}\sum_{i=1}^{n-1}JSD(R_{i},R_{i+1})$ (7)
where $n$ represents the number of time steps in the video data and $R_{i}$ represents the video representations at the $i$-th time steps.
Constrained representations can be regarded as temporal-invariant representations. Our proposed model, therefore, not only focuses on the similarity between frames but also captures global consistency features across the time sequence through temporal-invariant learning.
II-C Text-Driven Fusion Module
Fusion Procedure. Text-Driven Fusion Module has two parallel inter-modality attention streams with respective gate-controlled mechanisms. To enhance modality alignment, both streams are driven by the text modality to provide consistent context.
After positional encoding, the modality-specific features $S_{i}$ ($i\in\{a,v\}$) and $S_{t}$ are processed through the cross-attention stream to produce the interacted features $F_{ti}$ ($i\in\{a,v\}$):
$F_{ti}=Attention(S_{t},S_{i},S_{i})=softmax\left(\frac{S_{t}S_{i}^{T}}{\sqrt{d_{k}}}\right)S_{i}$ (8)
Meanwhile, the gated mechanism takes modality-invariant features $I_{i}$ ($i\in\{a,v\}$) and $I_{t}$ as inputs, producing strong correlations between modality-invariant features at each time step to control the interactions of modality-specific features.
Different from the previous work [12], we use modality-invariant representations to guide the interactions, rather than fused modality-specific representations themselves. During modality representation learning, modality-specific features may develop more distinct representations, which can make it challenging to accurately assess the similarity between modalities during fusion.
Since the modality-invariant features capture the common information across different modalities, we believe that using the modality-invariant features provides a more reliable basis for assessing the correlation between different modalities features. Our gated mechanism employs the Factorized Bilinear Pooling (FBP) [13] to generate the temporal gated signals $S_{g}$ ($g\in\{a,v\}$), as illustrated in Fig. 1. The formulaic expression can be given as:
$F_{mul}=(S_{t}W_{Q})\cdot(S_{i}W_{K})$ (9)
$F_{sp}=SumPool(F_{mul},k)$ (10)
$F_{norm}=F_{sp}/\|F_{sp}\|_{2}$ (11)
$S_{g}=F_{norm}W_{norm}$ (12)
The final representation $F_{final}$ is defined as:
$F_{final}=concatenate(S_{a}\cdot F_{ta},S_{v}\cdot F_{tv})$ (13)
Prediction. We feed the fused representation into an MLP to obtain the prediction output. The final loss function is expressed as follows:
$\mathcal{L}=\mathcal{L}_{task}+\alpha\mathcal{L}_{con}+\beta\mathcal{L}_{ti}++\gamma\mathcal{L}_{dom}$ (14)
where $\alpha$, $\beta$, and $\gamma$ are the trade-off parameters and $L_{task}\in\{L_{MSE},L_{CE}\}$ stands for the loss prediction function for different tasks.
TABLE I: The Experiment Results on CMU-MOSI and CMU-MOSEI
Models CMU-MOSI CMU-MOSEI
MAE Corr Acc-2 F1-Score Acc-7 MAE Corr Acc-2 F1-Score Acc-7
MISA 0.783 0.761 81.8/83.4 81.7/83.6 42.3 0.555 0.756 83.6/85.5 83.8/85.3 52.2
RegBn - 0.691 81.8/- 82.3/- 38.6 - 0.666 81.1/- 81.2/- 50.5
MMIN 0.741 0.795 83.53/85.52 83.46/85.51 - 0.542 0.761 83.84/85.88 83.91/85.76 -
ConFEDE 0.742 0.784 84.17/85.52 84.13/85.52 42.27 0.522 0.780 81.65/85.82 82.17/85.83 54.86
CAGC 0.775 0.774 -/85.70 -/85.60 44.80 - - - - -
Self-MM 0.713 0.798 84.00/85.98 84.42/85.95 - 0.530 0.765 82.81/85.17 82.53/85.30 -
FDMER 0.724 0.788 84.6/- 84.7/- 44.1 0.536 0.773 86.1/- 85.8/- 54.1
RTIL 0.683 0.814 85.13/86.89 85.08/86.90 45.63 0.528 0.795 86.12/86.55 85.97/86.21 52.56
a The best results are labeled in bold.
III EXPERIMENTS
III-A Datasets
In this paper, we validate our approach on two widely used multimodal sentiment recognition datasets, which are CMU-MOSI [14] and CMU-MOSEI [15]. CMU-MOSI contains 2,199 opinion segments. Each sample is annotated with a sentiment score on the scale ranging from negative to positive [-3, 3]. CMU-MOSEI comprises 23,453 annotated video clips from 1,000 speakers, each annotated with a sentiment scale from -3 to 3. We use the segmentation methods provided by the CMU-Multimodal SDK [16] in our experiments.
III-B Evaluation Criteria
Following the previous works [6, 19, 20], we utilize five metrics to evaluate the performance of the proposed model. For the sentiment classification task, we report binary classification accuracy (Acc-2), seven-class classification accuracy (Acc-7) and weighted F1 score (F1-Score). For the regression task, we report mean absolute error (MAE) and Pearson correlation (Corr). For Acc-2 and F1-Score, we use the segmentation marker -/- to report the results, where the left score corresponds to ”negative/non-negative” classification, and the right score corresponds to ”negative/positive” classification.
III-C Comparison with Baselines
To evaluate the rationality and effectiveness of our method, we compare the proposed model with the following recent and competitive baselines: MISA [6], RegBn [17], CAGC [18], Self-MM [19], ConFEDE [20], and MMIN [21].
The results compared with baselines on the two datasets are presented in TABLE I. We have the following observations. Our method significantly outperforms the previous state-of-the-art methods across all metrics on both benchmarks except for the seven-class classification task and MAE on the CMU-MOSEI dataset. The reason for the performance degradation in the seven-class classification task and MAE is that the model overly focuses on the adversarial learning task rather than the inter-class classification task.
Compared with the MISA [6], which learns different subspace representations as well, our method demonstrates that using an adversarial manner can better disentangle modality-invariant and modality-specific subspaces. Compared with the recent MMIN [21], witch exploits the unique characteristics of different modalities at a coarse-grained level, our text-driven modality fusion approach effective learns multimodal representations with a simpler structure.
III-D Ablation Studies
We conducted several modality absence studies to evaluate the modality robustness of our method. As shown in TABLE II, when one modality is absent, the performance of models slightly decreases, but the overall accuracy remains competitive.
To evaluate the noise robustness, we add some Gaussian noise to the initially extracted features, following the N(0, 0.5) distribution. TABLE III demonstrates that the addition of noise does not significantly impact the model performance, and some metrics even show a slight improvement. Compared with the MISA [6], our model exhibits less performance degradation.
TABLE II: The Modality Ablation Study
A T V Acc-2 F1 MAE Corr Acc-7
78.28/81.55 77.27/82.26 0.83 0.80 35.42
83.97/86.43 83.88/86.44 0.696 0.805 43.44
83.38/85.82 83.22/85.90 0.692 0.806 44.17
85.13/86.89 85.07/86.90 0.683 0.814 45.63
TABLE III: The Noise Robustness Experiments
Models Noise Acc-2 F1 MAE Corr Acc-7
RTIL 85.13/86.89 85.07/86.90 0.683 0.814 45.63
84.99/86.89 84.91/86.91 0.681 0.814 45.63
MISA* 80.90/82.93 80.86/82.95 0.809 0.753 42.27
80.17/82.32 80.11/82.30 0.807 0.754 41.11
aMISA with $\ast$ are reproduced under the same conditions.
IV Conclusion
In this paper, we present a novel multimodal sentiment recognition model to effectively learn representations from different modalities. To enhance temporal consistency of modality representations, we introduce the concept of temporal-invariant learning for the first time. Due to the superiority of the text modality, we design a text-driven fusion model, gated by modality-invariant representations adaptively. Furthermore, adversarial learning facilitates the disentanglement of the modality representation space. Experimental results demonstrate the superiority of our approach in multimodal sentiment recognition tasks.